Question 9 Not complete Points out of 1.00 P Flag question During a workout, the football players ran up the stadium stairs in 89.64 seconds. The stairs are 130 m long and inclined at an angle of 30 degrees. If a typical player weighs 191 pounds, estimate the average power output of the player in hP, on the way up. Ignore friction and air resistance. (1 kg weighs 2.2 lbs, 746 W 1 hP) Answer: Check
Answer #1

here,

l = 130 m

theta = 30 degree

the height of stairs , h = l * sin(theta) = 65 m

weight , W = 191 pounds

mass , m = 86.6 kg

the average power output , P = potential energy gained /time

P = m * g * h /( t)

P = 86.6 * 9.81 * 65 /89.64

P = 616.025 W

P = 0.83 hp